OpenStudy (anonymous):

What is the general rule for finding the x-intercept of a function? Is it to factor and set equal to zero?

OpenStudy (anonymous):

For example \[\left| 1divx \right|\]

OpenStudy (anonymous):

lemme try that again... f(x) = \[|1/x |\]

OpenStudy (anonymous):

sorry toally looking at the wong problem For ex. f(x) = \[x ^{2} - x -2\]

OpenStudy (radar):

The x intercepts are points where y=0 (point is on x axis. y f(x)=x^2-x-2 0=x^2-x-2 or as you suggested factor, (x-2)(x+1)=0 if this is true then one or both of the factors equal 0. This results in x=2 and x=-2

OpenStudy (anonymous):

thanks!